6x-3y=18, 6x+3y=-12 Using the Elimination Method, please show all work.
Answer by richwmiller(17219)

add them
12x=6
x=1/2
plug in 1/2 for x
6*1/2-3y=18
3-3y=18
divide by 3
1-y=6
subtract 6
-5-y=0
add y
-5=y